suppose y varies inversly as x and y=24 when x=3.find x when y=36.
Answer by stanbon(75887)

suppose y varies inversely as x and y=24 when x=3.find x when y=36.
--------------------------------------
y = k/x expresses the inverse relationship.
----------------------------
Solve for "k":
24 = k/3
k = 72
So the equation is y = 72/x
---------------------------------------
Now find x when y = 36:
36 = 72/x
x = 72/36
x = 2
=========================================
Cheers,
Stan H.